Four times the difference of a number and 3 is the same as 3 increased by four times the number plus 3 times the number. Find the number
Answer by mananth(16949)

number = x
difference of x & 3 = x-3
4 times
4(x-3)
..
4 times the number= 4x
3+4x +3x
3+7x
4(x-3)=3+7x
4x-12=3+7x
3x=-15
/3
x=-5